In 2004, three wide receivers for the Indianapolis Colts caught a total of 37 touchdown passes. Reggie Wayne caught 2 more then Brandon Stokely and Marvin Harrison caught 3 more than Reggie Wayne. How many touchdown passes did each receiver catch?
Answer by stanbon(75887)

In 2004, three wide receivers for the Indianapolis Colts caught a total of 37 touchdown passes. Reggie Wayne caught 2 more then Brandon Stokely and Marvin Harrison caught 3 more than Reggie Wayne. How many touchdown passes did each receiver catch?
---------------------
Equations:
W + S + H = 37
W = S + 2
H = W + 3
----------
Substitute and solve for W
W + W-2 + W+3 = 37
----
3W = 36
W = 12 (# caught by Wayne)
S = W-2 = 10 (# caught by Stokely)
H = W+3 = 15 (# caught by Harrison)
=========================================
Cheers,
Stan H.
===================